A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1880–1889

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

17 September 2022 Version 1.02

Title word cross-reference

14th [JGB84]. 1883 [Ano84c, Ano84b, Gar83a].

25th [Gar83a].

Acetic [Cla87]. Acids [Cla87]. Act [Ano87a]. Affect [Che85]. Afternoon [Ano84a, Ano86a]. Agent [Ric84]. Alaska [Mur84]. Albatross [Was86]. Amended [Ano84b]. Amendments [Ano84c]. America [Bea84, Gar83b]. American [Ano84c, Ano84i, Ano85a, Ano86d, Ano87e, Ano88e, Ano89b, Ano89a, Ano89c, Ano80b, Ano81a, Ano82b, Ano83b, Ano84e, Ano85b, Ano86e, Ano87c, Ano88a, Ano89d, Atw84, Col84, Gil85]. Annual [Ano80c, Ano81d, Ano82a, Ano83e, Ano84b, Ano84i, Ano85a, Ano86d, Ano87e, Ano88e, Ano89b, Ano89a, Ano89c]. Arctic [Mur84]. Art [Bis86]. Artificial [Sto84]. Association [Ano83e, Ano84c, Ano80b, Ano80c, Ano81a, Ano81d, Ano82a, Ano82b, Ano83b, Ano84e, Ano84i, Ano85a]. Atlantic [Bla83]. Automatic [McD83].

Barrow [Mur84]. Basin [Sto84]. Bass [Gre82, Hen83, Hen85, Hol88, Pag84, Wor84]. Bedford [Gar83a]. Beds [Bla85]. Beginning [Gar83b]. Biennial [Atk85]. Billingsgate [Cox85]. Black [Hen83, Hen85, Hol88, Pag84]. Boracic [Cla87]. Both [Mat81]. Breeding [Ryd82]. British [Cox86]. Bucksport [Atk85]. Building [M’D83].

can [Mat81]. Cape [Tru85b]. Carolina [Wor87]. Carp [M’G82]. Cause [Ryd82]. Causes [Mc’84]. Central [Ano86b]. Certain [Che85]. Changes [Atw87, Smi82, Ano84c]. Characteristics [Gil85]. Chemical [Atw81, Atw84, Atw87]. Chief [Gil85]. Claims [Ste85]. Co [Bis88]. Co-operation [Bis88]. Coast [Bla83]. Cod [Col84]. Cod-Fisheries [Col84]. Cold [Mat84, Mat85b, Mat86c, Mat87]. Color [Goo89, Ryd82]. Columbia [Sto84]. Commission [Nev88]. Commissioners [Ear87]. Committee [Ano88d, Ano80d, Ano81b, Ano82c, Ano83c, Ano84f, Ano85c, Ano86f, Ano87d, Ano88c, Ano89e]. Comparative [Hen84]. Composition [Atw81, Atw84]. Compounds [Cla87]. Condition [Win84]. Connecticut [Hud84]. Connection [Ryd85]. Constitution [Ano80a, Ano86c, Ano87b, MCC85, Ano84c, Ano84b]. Contrivances [Ryd85]. Crustaceans [Mat83b]. Cultural [Ano80b, Ano80c, Ano81a, Ano81d, Ano82a, Ano82b, Ano83b, Ano83e, Ano84c, Ano84e, Ano84i, Ano85a, Ano86b]. Culture [Ann84, Bis86, Gar83b, Goo81b, Joh83, M’D85, Mat86a, Wor87, Bis88, O’B88].

Day [Ano80f, Ano81c, Ano82d, Ano86g, Ano83e, Ano86d, Ano89b, Ano89c]. Decade [Smi82]. Decrease [Rat84]. Deep [Was86]. Destruction [Ric84]. Determine [Ryd84]. Developed [Ryd85]. Development [Mat82, McD83]. Digestibility [Atw88]. Distribution [Hen83, Jor88, Ste86]. Does [Che85]. Dredging [Was86]. During [Smi82].

Eel [Goo81a]. Effect [Atw87]. Eighteenth [Ano89b, Ano89a, Ano89c]. Eleventh [Ano82a]. Embryo [Mat82]. Embryos [Ryd84]. Encouragement [Wor87]. Endurance [M’G82]. England [GC82]. Epochs [Goo81b]. Erie [Cla85]. Esquimaux [Ano85f]. Evening [Ano89a]. Excellence [Hen84]. Excursion [Ano84g, Ano85d]. Executive [Ano80d, Ano81b, Ano82c, Ano83c, Ano84f, Ano85c, Ano86f, Ano87d, Ano88c, Ano89e]. Experiments [Atk85, McD83, Par88a, Ric83]. Extermination [Mat83a]. Eyed [Nev87].

Facts [Bla83]. Fauna [Gil85]. Few [Bla83]. Fifteenth [Ano86d]. First [Ano83e, Ano86d, Ano89b]. Fish [Ann84, Ano80b, Ano80c, Ano80e, Ano81a, Ano81d, Ano82a, Ano82b, Ano83b, Ano83e, Ano84c, Ano84e, Ano84i, Ano85a, Ano86b, Atw81, Atw88, Bis86, Bis88, Che83, Cla86, Cla87, Cox86, Ear87, Gar83b, Gil85, Goo81b, M’D83, M’D85, McD83, Mc’84, Mur84, Nev88, O’B88, Ric84, Ryd84, Tom87]. Fish-Cultural [Ano82a, Ano82b, Ano83b, Ano83e, Ano84c, Ano84i, Ano85a, Ano86b]. Fish-Culture [Ann84, Bis86, M’D85, Bis88, O’B88]. Fish-Hatching [McD83]. Fish-Way [M’D83]. Fisheries [Ano85a, Ano85b, Ano86d, Ano86e, Ano87c, Ano87e, Ano88a, Ano88e, Ano89b, Ano89a, Ano89c, Ano89d, Bla84, Col84, Hal80, Hud84, Ito87, Smi82]. Fishery [GC82, Tru85b, Wil84]. Fishes [Atw84, Bea84, Bla83, Che85, For88, Goo89, Gre81, Hen84, Jor88, Mat81, Ryd85]. Fishing [Mur84]. Floating [Atw87]. Florida [Wil84]. following [Ano84h]. Food [Atw84, Bla83, Che83, Che85, For88, Hen84, O’B88, Ryd82]. Forces [Ryd84]. Fourteenth [Ano85a]. Fresh [Jor88, Mat81, Mat84]. Fresh-water [Jor88]. Frozen [Cla86]. Fry [Par88a]. Fulton [Ano80e]. Fungus [Ric84]. Future [Win84].

Game [Che85]. General [Phi81]. Giant [Ste85]. Gill [Col84]. Gill-Nets [Col84]. Glance [Cox85]. Grayling [Par88b, Whi86]. Great [Smi82]. Green [Ryd82]. Growth [M’G82].

Habits [M’G82, Mat83a]. Haddock [GC82]. Harbor [Mat86c, Mat84, Mat85b, Mat87]. Hatching [Gre82, Mat84, Mat85a, Mat86b, McD83, Nev87]. Hatteras [Tru85b]. Head [Ryd89]. Hibernation [Hen85]. History [Cla86, Goo81b, Goo82, McD83]. Hudson [Dan87, Mat89]. Hyaline [Ryd89]. Hybridizing [Gre81]. Hybrids [Bea89].

Iced [Cla86]. Including [Ano84b]. Index [Ano84d]. Industry [Goo84, Win84]. Influencing [Mc’84]. Intentional [Ste86]. Introduction [Col84]. Invertebrates [Atw84]. Isles [Cox86].

Japan [Ito87]. Jar [McD83].

Labrador [Hal80]. Lake [Cla85, Tom87]. Lakes [Smi82]. Land [Atk84]. Land-Locked [Atk84]. last [Ano84c]. Lateral [Ryd89]. Leading [McD83]. Legislation [Bla84]. letter [Ano84h]. Line [Ryd89]. Live [Mat81]. Lobster [Joh83]. Lobsters [Ano83a, Rat84]. Locked [Atk84].

Maine [Pag84]. Market [Ano80e]. Materials [Goo82]. May [Gar83a, JGB84]. Meeting [Ano80c, Ano81d, Ano82a, Ano83e, Ano84c, Ano84b, Ano84i, Ano85a, Ano86d, Ano87e, Ano88e, Ano89b, Ano89a, Ano89c]. Members [Ano80b, Ano81a, Ano82b, Ano83b, Ano84e, Ano85b, Ano86e, Ano87c, Ano88a, Ano89d]. Michigan [But89, Whi86]. Migration [Tom87].

Mississippi [For88]. Movements [Mc’84].

N.Y. [Mat86c]. Natural [Mc’84, O’B88]. Necessary [Bla84]. Nets [Col84]. Ninth [Ano80c]. North [Bea84, Gil85, Wor87]. Notes [Ann84, Ano88b, Atk84, But89, For88, Rat84, Ryd82]. Nutritive [Atw81, Atw84, Atw87].

O [Gar83a]. Objective [M’D85]. Observations [Hol88, Par88b]. Ocean [Bla84]. Office [Ano84h]. Officers [Ano80d, Ano81b, Ano82c, Ano83c, Ano84f, Ano85c, Ano86f, Ano87d, Ano88c, Ano89e]. operation [Bis88]. organization [Ano84c]. Organs [Ryd89]. other [Cla87]. Our [Atw84, Van85]. Ova [Ryd85]. Oyster [Bla85, Goo84, Mat86a, Ric83, Ryd82, Win84]. Oysters [Atw87].

Pertaining [Ann84]. Pike [Nev87]. Planting [Cla85]. Poachers [Ann81]. Point [Mur84]. Points [M’D85]. Porpoise [Tru85b, Tru85a]. Practical [Bis86]. Present [Win84]. Preservation [Cla87]. Proceedings [Ano80f]. Produced [Atw87]. Propagation [O’B88, Ric83, Sto84, Wor84]. Proposed [MCC85]. Prospects [Win84]. Protecting [Mat85a]. Protection [Bla84]. Protective [Ryd85]. Publication [Ano88d]. Puget [Ste85].

Qualities [Che85]. Question [Goo81a].

Rainbow [Ann82]. read [Ano84h]. Reference [O’B88]. Relation [Bla83]. Remarkable [Mat82]. Report [Ano80g, Ano81e, Ano82e, Ano83d, Ano84j, Ano85e, Ano86h, Ano87f, Ano88f, Ano89f]. Restore [Van85]. Result [Col84]. Results [Cla85]. River [Ano84g, Ano85d, Sto84, Mat89]. Rivers [Mc’84].

Sales [Ano80e]. Salicylic [Cla87]. Salmon [Atk84, Atk85, Dan87, Mat82, Mat89, Sto84]. Salmonidæ [Bea89]. Salt [Mat81, Mat84, Ric84]. Salt-Water [Mat84]. Sea [Was86]. Seasons [Bla83]. Second [Ano80f, Ano81c, Ano82d, Ano86g, Ano89c]. Secretary [Ano84h]. Session [Ano84a, Ano86a, Ano89a]. Seventeenth [Ano88e]. Shad [Ryd89]. Shell [Hud84]. Shellfish [Wor87]. Shore [Hal80]. Shore-Fisheries [Hal80]. Sixteenth [Ano87e]. Smelt [Mat85a, Mat86b]. Society [Ano85a, Ano86b, Ano86d, Ano89b, Ano89a, Ano89c, Ano85b, Ano86e, Ano87c, Ano87e, Ano88a, Ano88e, Ano89d]. Some [Hol88, M’D85, Par88a, Par88b, Ryd85]. Sound [Ste85]. Spawning [Atk85, Bla83]. Special [O’B88]. Species [Ryd85, Ste86]. Sponge [Wil84]. Spring [Mat84, Mat85b, Mat86c, Mat87]. State [Ear87]. States [Cla86]. Statistics [Phi81]. Steak [Tru85a]. Steamer [Was86]. Stick [Ano85f]. Streams [Van85]. Striped [Gre82, Wor84]. Sturgeon [Gre82]. Sunfish [Mat83a]. Superior [Tom87]. Survival [Ryd84]. Swordfish [Goo82]. System [M’D83].

Tenth [Ano81d]. Their [Mat83a, Atw87]. Thirteenth [Ano84i]. Throwing [Ano85f]. Tissues [Ryd89]. Trade [Cla86]. Transplanting [Che85]. Transportation [Mat83b]. Transporting [Cox86]. Treasurer [Ano80g, Ano81e, Ano82e, Ano83d, Ano84j, Ano85e, Ano86h, Ano87f, Ano88f, Ano89f]. Trout [Ann82, But89, Gre82, Van85]. Twelfth [Ano84b, Ano83e].

U.S. [Was86]. Unintentional [Ste86]. United [Cla86]. upon [Atw87, Par88b]. Use [Ano85f, Cla87].

Valley [For88]. Value [Atw81, Atw84, Atw87]. Various [Ryd85].

Wall [Nev87]. Wall-Eyed [Nev87]. Washington [Ano84h]. Water [Mat81, Mat84, Jor88]. Way [M’D83]. Wednesday [JGB84]. Which [Mat81, Ryd84]. White [Bea84]. Whitefish [Cla85, Par88a]. Winter [GC82]. Wisconsin [Nev88]. Work [But89, Mat85b, Mat86c, Mat87, Nev88]. World [Goo84].

York [Bla85].

References

Annin:1881:P

[Ann81] James Annin, Jr. Poachers. Transactions of the American Fisheries Society, 10(1):76–81, January 1881. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Annin:1882:RT

[Ann82] James Annin, Jr. The rainbow trout. Transactions of the American Fisheries Society, 11(1):20–24, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Annin:1884:NPF

[Ann84] James Annin, Jr. Notes pertaining to fish-culture. Transactions of the American Fisheries Society, 13(1):109–111, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1880:C

[Ano80a] Anonymous. Constitution. Transactions of the American Fisheries Society, 9(1):66, January 1880. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1880:MAF

[Ano80b] Anonymous. Members of the American Fish Cultural Association. Transactions of the American Fisheries Society, 9(1):68–73, January 1880. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1880:NAM

[Ano80c] Anonymous. Ninth Annual Meeting of The Fish Cultural Association. Transactions of the American Fisheries Society, 9(1):3–33, January 1880. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1880:OEC

[Ano80d] Anonymous. Officers, 1880–1 and Executive Committee. Transactions of the American Fisheries Society, 9(1):2, January 1880. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1880:SFF

[Ano80e] Anonymous. Sales of fish in Fulton market. Transactions of the American Fisheries Society, 9(1):43–44, January 1880. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1880:SDP

[Ano80f] Anonymous. Second day’s proceedings. Transactions of the American Fisheries Society, 9(1):44–65, January 1880. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1880:TR

[Ano80g] Anonymous. Treasurer’s report. Transactions of the American Fisheries Society, 9(1):67, January 1880. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1881:MAF

[Ano81a] Anonymous. Members of the American Fish Cultural Association. Transactions of the American Fisheries Society, 10(1):133–136, January 1881. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1881:OEC

[Ano81b] Anonymous. Officers, 1881–82 and Executive Committee. Transactions of the American Fisheries Society, 10(1):2, January 1881. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1881:SD

[Ano81c] Anonymous. Second day. Transactions of the American Fisheries Society, 10(1):60, January 1881. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1881:TAM

[Ano81d] Anonymous. Tenth Annual Meeting of the Fish Cultural Association. Transactions of the American Fisheries Society, 10(1):3–5, January 1881. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1881:TR

[Ano81e] Anonymous. Treasurer’s report. Transactions of the American Fisheries Society, 10(1):132, January 1881. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1882:EAM

[Ano82a] Anonymous. Eleventh Annual Meeting of the Fish-Cultural Association. Transactions of the American Fisheries Society, 11(1):3–5, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1882:MAF

[Ano82b] Anonymous. Members of the American Fish-Cultural Association. Transactions of the American Fisheries Society, 11(1):153–157, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1882:OEC

[Ano82c] Anonymous. Officers, 1882–83 and Executive Committee. Transactions of the American Fisheries Society, 11(1):2, January 1882. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1882:SD

[Ano82d] Anonymous. Second day. Transactions of the American Fisheries Society, 11(1):25–27, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1882:TR

[Ano82e] Anonymous. Treasurer’s report. Transactions of the American Fisheries Society, 11(1):151, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1883:L

[Ano83a] Anonymous. Lobsters. Transactions of the American Fisheries Society, 12(1):13–18, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1883:MAF

[Ano83b] Anonymous. Members of the American Fish-Cultural Association. Transactions of the American Fisheries Society, 12(1):79–83, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1883:OEC

[Ano83c] Anonymous. Officers, 1883–84 and Executive Committee. Transactions of the American Fisheries Society, 12(1):2, January 1883. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1883:TR

[Ano83d] Anonymous. Treasurer’s report. Transactions of the American Fisheries Society, 12(1):77, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1883:TAM

[Ano83e] Anonymous. Twelfth Annual Meeting of the Fish-Cultural Association: First Day. Transactions of the American Fisheries Society, 12 (1):3–5, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1884:AS

[Ano84a] Anonymous. The afternoon session. Transactions of the American Fisheries Society, 13(1):170–171, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1884:CAI

[Ano84b] Anonymous. Constitution: As Amended up to and Including the Twelfth Annual Meeting in 1883. Transactions of the American Fisheries Society, 13(1):vi–vii, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1884:CAF

[Ano84c] Anonymous. Constitution of the American Fish-Cultural Association: With all its Amendments and changes from its organization to its last Meeting in 1883. Transactions of the American Fisheries Society, 13 (1):iii–vi, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1884:I

[Ano84d] Anonymous. Index. Transactions of the American Fisheries Society, 13(1):ix, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1884:MAF

[Ano84e] Anonymous. Members of the American Fish Cultural Association. Transactions of the American Fisheries Society, 13(1):249–253, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1884:OEC

[Ano84f] Anonymous. Officers, 1884–85 and Executive Committee. Transactions of the American Fisheries Society, 13(1):ii, January 1884. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1884:RE

[Ano84g] Anonymous. The river excursion. Transactions of the American Fisheries Society, 13(1):230–247, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1884:STR

[Ano84h] Anonymous. The secretary then read the following letter: Washington Office. Transactions of the American Fisheries Society, 13(1):229–230, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1884:TAM

[Ano84i] Anonymous. Thirteenth Annual Meeting of the American Fish-Cultural Association. Transactions of the American Fisheries Society, 13(1):5, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1884:TR

[Ano84j] Anonymous. Treasurer’s report. Transactions of the American Fisheries Society, 13(1):248, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1885:FAM

[Ano85a] Anonymous. Fourteenth Annual Meeting of the American Fisheries Society: American Fish-Cultural Association. Transactions of the American Fisheries Society, 14(1):5–8, January 1885. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1885:MAF

[Ano85b] Anonymous. Members of the American Fisheries Society. Transactions of the American Fisheries Society, 14(1):101–106, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1885:OEC

[Ano85c] Anonymous. Officers for 1885–86 and Executive Committee. Transactions of the American Fisheries Society, 14(1):2, January 1885. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1885:RE

[Ano85d] Anonymous. The river excursion. Transactions of the American Fisheries Society, 14(1):97–98, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1885:TR

[Ano85e] Anonymous. Treasurer’s report. Transactions of the American Fisheries Society, 14(1):99, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1885:UTS

[Ano85f] Anonymous. The use of the throwing stick by the Esquimaux. Transactions of the American Fisheries Society, 14(1):66–68, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1886:AS

[Ano86a] Anonymous. Afternoon session. Transactions of the American Fisheries Society, 15(1):9–10, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1886:CFC

[Ano86b] Anonymous. The central fish-cultural society. Transactions of the American Fisheries Society, 15(1):25, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1886:C

[Ano86c] Anonymous. Constitution. Transactions of the American Fisheries Society, 15(1):3–4, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1886:FAM

[Ano86d] Anonymous. Fifteenth Annual Meeting of the American Fisheries Society: First Day. Transactions of the American Fisheries Society, 15(1):5–9, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1886:MAF

[Ano86e] Anonymous. Members of the American Fisheries Society. Transactions of the American Fisheries Society, 15(1):95–100, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1886:OEC

[Ano86f] Anonymous. Officers for 1886–7 and Executive Committee. Transactions of the American Fisheries Society, 15(1):2, January 1886. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1886:SD

[Ano86g] Anonymous. Second day. Transactions of the American Fisheries Society, 15(1):22–24, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1886:TR

[Ano86h] Anonymous. Treasurer’s report. Transactions of the American Fisheries Society, 15(1):93, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1887:A

[Ano87a] Anonymous. The Act. Transactions of the American Fisheries Society, 16(1):54–59, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1887:C

[Ano87b] Anonymous. Constitution. Transactions of the American Fisheries Society, 16(1):3–4, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1887:MAF

[Ano87c] Anonymous. Members of the American Fisheries Society. Transactions of the American Fisheries Society, 16(1):67–72, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1887:OEC

[Ano87d] Anonymous. Officers for 1887–8 and Executive Committee. Transactions of the American Fisheries Society, 16(1):2, January 1887. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1887:SAM

[Ano87e] Anonymous. Sixteenth Annual Meeting of the American Fisheries Society. Transactions of the American Fisheries Society, 16(1):5–8, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1887:TR

[Ano87f] Anonymous. Treasurer’s report. Transactions of the American Fisheries Society, 16(1):65–66, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1888:MAF

[Ano88a] Anonymous. Members of the American Fisheries Society. Transactions of the American Fisheries Society, 17(1):110–115, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1888:N

[Ano88b] Anonymous. Notes. Transactions of the American Fisheries Society, 17(1):104–107, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1888:OEC

[Ano88c] Anonymous. Officers for 1888–9 and Executive Committee. Transactions of the American Fisheries Society, 17(1):2, January 1888. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1888:PC

[Ano88d] Anonymous. The publication committee. Transactions of the American Fisheries Society, 17(1):107–108, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1888:SAM

[Ano88e] Anonymous. Seventeenth Annual Meeting of the American Fisheries Society. Transactions of the American Fisheries Society, 17(1):3–4, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1888:TR

[Ano88f] Anonymous. Treasurer’s report. Transactions of the American Fisheries Society, 17(1):109, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1889:EAMb

[Ano89a] Anonymous. Eighteenth Annual Meeting of the American Fisheries Society: Evening Session. Transactions of the American Fisheries Society, 18(1):31–36, January 1889. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1889:EAMa

[Ano89b] Anonymous. Eighteenth Annual Meeting of the American Fisheries Society: First Day. Transactions of the American Fisheries Society, 18(1):3–11, January 1889. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1889:EAMc

[Ano89c] Anonymous. Eighteenth Annual Meeting of the American Fisheries Society: Second Day. Transactions of the American Fisheries Society, 18(1):37–38, January 1889. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1889:MAF

[Ano89d] Anonymous. Members of the American Fisheries Society. Transactions of the American Fisheries Society, 18(1):80–87, January 1889. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1889:OEC

[Ano89e] Anonymous. Officers for 1889–90 and Executive Committee. Transactions of the American Fisheries Society, 18(1):2, January 1889. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1889:TR

[Ano89f] Anonymous. Treasurer’s report. Transactions of the American Fisheries Society, 18(1):79, January 1889. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1884:NLL

[Atk84] Charles G. Atkins. Notes on land-locked salmon. Transactions of the American Fisheries Society, 13(1):40–56, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atkins:1885:BSS

[Atk85] Charles G. Atkins. The biennial spawning of salmon: (the Bucksport experiments). Transactions of the American Fisheries Society, 14(1): 89–94, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atwater:1881:CCN

[Atw81] W. O. Atwater. The chemical composition and nutritive value of fish. Transactions of the American Fisheries Society, 10(1):124–131, January 1881. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atwater:1884:CCN

[Atw84] W. O. Atwater. The chemical composition and nutritive value of our American food fishes and invertebrates. Transactions of the American Fisheries Society, 13(1):171–194, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atwater:1887:CCP

[Atw87] W. O. Atwater. The chemical changes produced in oysters in floating, and their effect upon the nutritive value. Transactions of the American Fisheries Society, 16(1):37–52, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atwater:1888:DF

[Atw88] W. O. Atwater. The digestibility of fish. Transactions of the American Fisheries Society, 17(1):69–83, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bean:1884:WFN

[Bea84] Tarleton H. Bean. The white fishes of North America. Transactions of the American Fisheries Society, 13(1):32–39, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bean:1889:HS

[Bea89] Tarleton H. Bean. Hybrids in Salmonidæ. Transactions of the American Fisheries Society, 18(1):12–20, January 1889. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bissell:1886:FCP

[Bis86] John H. Bissell. Fish-culture — a practical art. Transactions of the American Fisheries Society, 15(1):36–50, January 1886. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bissell:1888:COF

[Bis88] John H. Bissell. Co-operation in fish-culture. Transactions of the American Fisheries Society, 17(1):89–100, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blackford:1883:FFR

[Bla83] E. G. Blackford. A few facts in relation to the food and spawning seasons of fishes on the Atlantic Coast. Transactions of the American Fisheries Society, 12(1):5–10, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blackford:1884:LNP

[Bla84] Eugene G. Blackford. Is legislation necessary for the protection of the ocean fisheries? Transactions of the American Fisheries Society, 13 (1):60–66, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blackford:1885:OBN

[Bla85] Eugene G. Blackford. The oyster beds of New York. Transactions of the American Fisheries Society, 14(1):85–89, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butler:1889:NTW

[But89] Wm A. Butler, Jr. Notes on trout work in Michigan. Transactions of the American Fisheries Society, 18(1):25–31, January 1889. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cheney:1883:FFF

[Che83] A. N. Cheney. Food fish and fish food. Transactions of the American Fisheries Society, 12(1):27–33, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cheney:1885:DTA

[Che85] A. N. Cheney. Does transplanting affect the food or game qualities of certain fishes? Transactions of the American Fisheries Society, 14 (1):55–59, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1885:RPW

[Cla85] Frank N. Clark. Results of planting whitefish in Lake Erie. Transactions of the American Fisheries Society, 14(1):40–50, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1886:HIF

[Cla86] A. Howard Clark. History of the iced fish and frozen fish trade of the United States. Transactions of the American Fisheries Society, 15 (1):68–83, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1887:FPU

[Cla87] A. Howard Clark. Fish preservation by the use of acetic, boracic, salicylic, and other acids and compounds. Transactions of the American Fisheries Society, 16(1):28–36, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:1884:RIG

[Col84] J. W. Collins. Result of the introduction of gill-nets into the American cod-fisheries. Transactions of the American Fisheries Society, 13(1): 212–228, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cox:1885:GB

[Cox85] William Van Zandt Cox. A glance at Billingsgate. Transactions of the American Fisheries Society, 14(1):76–85, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cox:1886:TFB

[Cox86] W. V. Cox. Transporting fish in the British Isles. Transactions of the American Fisheries Society, 15(1):56–58, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Danaher:1887:SH

[Dan87] F. M. Danaher. Salmon in the Hudson. Transactions of the American Fisheries Society, 16(1):59, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Earll:1887:SFC

[Ear87] R. Edward Earll. State Fish Commissioners. Transactions of the American Fisheries Society, 16(1):23–28, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Forbes:1888:NFF

[For88] S. A. Forbes. Notes on the food of the fishes of the Mississippi Valley. Transactions of the American Fisheries Society, 17(1):37–67, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garlick:1883:BM

[Gar83a] T. Garlick. Bedford, O., May 25th, 1883. Transactions of the American Fisheries Society, 12(1):48, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garlick:1883:BFC

[Gar83b] T. Garlick. The beginning of fish culture in America. Transactions of the American Fisheries Society, 12(1):47, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goode:1882:WHF

[GC82] G. Brown Goode and J. W. Collins. The winter haddock fishery of New England. Transactions of the American Fisheries Society, 11(1):43–56, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gill:1885:CCN

[Gil85] Theodore Gill. The chief characteristics of the North American fish fauna. Transactions of the American Fisheries Society, 14(1):69–72, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goode:1881:EQ

[Goo81a] G. Brown Goode. The eel question. Transactions of the American Fisheries Society, 10(1):81–124, January 1881. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goode:1881:EHF

[Goo81b] G. Brown Goode. Epochs in the history of fish culture. Transactions of the American Fisheries Society, 10(1):34–59, January 1881. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goode:1882:MHS

[Goo82] G. Brown Goode. Materials for a history of the swordfish. Transactions of the American Fisheries Society, 11(1):84–150, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goode:1884:OIW

[Goo84] G. Brown Goode. The oyster industry of the world. Transactions of the American Fisheries Society, 13(1):146–148, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goode:1889:CF

[Goo89] G. Brown Goode. The color of fishes. Transactions of the American Fisheries Society, 18(1):65–78, January 1889. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Green:1881:HF

[Gre81] Seth Green. Hybridizing fishes. Transactions of the American Fisheries Society, 10(1):5–34, January 1881. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Green:1882:HSB

[Gre82] Seth Green. Hatching striped bass, sturgeon and trout. Transactions of the American Fisheries Society, 11(1):37–43, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hallock:1880:SFL

[Hal80] Charles Hallock. Shore-fisheries of Labrador. Transactions of the American Fisheries Society, 9(1):34–40, January 1880. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1883:DBB

[Hen83] J. A. Henshall. On the distribution of the black bass. Transactions of the American Fisheries Society, 12(1):21–27, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1884:CEF

[Hen84] James A. Henshall. Comparative excellence of food fishes. Transactions of the American Fisheries Society, 13(1):115–122, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henshall:1885:HBB

[Hen85] James A. Henshall. Hibernation of the black bass. Transactions of the American Fisheries Society, 14(1):12–17, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holt:1888:SOB

[Hol88] C. F. Holt. Some observations on the black bass. Transactions of the American Fisheries Society, 17(1):33–37, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hudson:1884:SFC

[Hud84] William M. Hudson. The shell fisheries of Connecticut. Transactions of the American Fisheries Society, 13(1):124–146, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ito:1887:FJ

[Ito87] K. Ito. The fisheries of Japan. Transactions of the American Fisheries Society, 16(1):17–23, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1884:WM

[JGB84] S. M. Johnson, T. Garlick, and C. J. Bottemanne. Wednesday, May 14th. Transactions of the American Fisheries Society, 13(1):122–124, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1883:LC

[Joh83] S. M. Johnson. Lobster culture. Transactions of the American Fisheries Society, 12(1):18–21, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jordan:1888:DFW

[Jor88] David Starr Jordan. The distribution of fresh-water fishes. Transactions of the American Fisheries Society, 17(1):4–29, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1881:FWC

[Mat81] Fred Mather. Fishes which can live in both salt and fresh water. Transactions of the American Fisheries Society, 10(1):65–75, January 1881. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1882:RDE

[Mat82] Fred Mather. Remarkable development of embryo salmon. Transactions of the American Fisheries Society, 11(1):7–20, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1883:STH

[Mat83a] Fred Mather. Sunfish: Their habits and extermination. Transactions of the American Fisheries Society, 12(1):10–13, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1883:TC

[Mat83b] Fred Mather. Transportation of crustaceans. Transactions of the American Fisheries Society, 12(1):46–47, January 1883. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1884:FSW

[Mat84] Fred Mather. Fresh and salt-water hatching at Cold Spring Harbor. Transactions of the American Fisheries Society, 13(1):6–15, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1885:PHS

[Mat85a] Fred Mather. Protecting and hatching the smelt. Transactions of the American Fisheries Society, 14(1):17–32, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1885:WCS

[Mat85b] Fred Mather. Work at Cold Spring Harbor. Transactions of the American Fisheries Society, 14(1):94–97, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1886:OC

[Mat86a] Fred Mather. Oyster culture. Transactions of the American Fisheries Society, 15(1):26–36, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1886:SH

[Mat86b] Fred Mather. Smelt hatching. Transactions of the American Fisheries Society, 15(1):10–16, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1886:WCS

[Mat86c] Fred Mather. Work at Cold Spring Harbor, N.Y. Transactions of the American Fisheries Society, 15(1):84–92, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1887:WCS

[Mat87] Fred Mather. Work at Cold Spring Harbor. Transactions of the American Fisheries Society, 16(1):8–13, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1889:SHR

[Mat89] Fred Mather. Salmon in the Hudson River. Transactions of the American Fisheries Society, 18(1):39–65, January 1889. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDonald:1884:NCI

[Mc’84] Marshall Mc’Donald. Natural causes influencing the movements of fish in rivers. Transactions of the American Fisheries Society, 13(1): 164–170, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mather:1885:PC

[MCC85] Fred Mather, W. V. Cox, and F. N. Clark. Proposed constitution. Transactions of the American Fisheries Society, 14(1):3–4, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDonald:1883:HEL

[McD83] Marshall McDonald. History of the experiments leading to the development of the automatic fish-hatching jar. Transactions of the American Fisheries Society, 12(1):34–46, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MDonald:1883:NSF

[M’D83] Marshall M’Donald. A new system of fish-way building. Transactions of the American Fisheries Society, 12(1):57–76, January 1883. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MDonald:1885:SOP

[M’D85] M. M’Donald. Some objective points in fish-culture. Transactions of the American Fisheries Society, 14(1):72–76, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

MGovern:1882:HEG

[M’G82] Hugh D. M’Govern. The habits, endurance, and growth of carp. Transactions of the American Fisheries Society, 11(1):5–7, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murdoch:1884:FFP

[Mur84] John Murdoch. Fish and fishing at Point Barrow, Arctic Alaska. Transactions of the American Fisheries Society, 13(1):111–115, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nevin:1887:HWE

[Nev87] James Nevin. Hatching the wall-eyed pike. Transactions of the American Fisheries Society, 16(1):14–16, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nevin:1888:WWF

[Nev88] Jas Nevin. Work of the Wisconsin Fish Commission. Transactions of the American Fisheries Society, 17(1):100–103, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OBrien:1888:PNF

[O’B88] M. E. O’Brien. The propagation of natural food for fish, with special reference to fish-culture. Transactions of the American Fisheries Society, 17(1):29–33, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Page:1884:BBM

[Pag84] George Shepard Page. Black bass in Maine. Transactions of the American Fisheries Society, 13(1):57–60, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:1888:SEF

[Par88a] J. C. Parker. Some experiments with the fry of whitefish. Transactions of the American Fisheries Society, 17(1):67–69, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parker:1888:SOU

[Par88b] J. C. Parker. Some observations upon the grayling. Transactions of the American Fisheries Society, 17(1):83–89, January 1888. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1881:GS

[Phi81] Barnet Phillips. General statistics. Transactions of the American Fisheries Society, 10(1):61–64, January 1881. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rathbun:1884:NDL

[Rat84] Richard Rathbun. Notes on the decrease of lobsters. Transactions of the American Fisheries Society, 13(1):201–208, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rice:1883:EOP

[Ric83] H. J. Rice. Experiments in oyster propagation. Transactions of the American Fisheries Society, 12(1):49–56, January 1883. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rice:1884:SAD

[Ric84] H. J. Rice. Salt as an agent for the destruction of the fish fungus. Transactions of the American Fisheries Society, 13(1):15–21, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ryder:1882:NBF

[Ryd82] John A. Ryder. Notes on the breeding, food and cause of the green color of the oyster. Transactions of the American Fisheries Society, 11 (1):57–84, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ryder:1884:FWD

[Ryd84] John A. Ryder. On the forces which determine the survival of fish embryos. Transactions of the American Fisheries Society, 13(1):195– 200, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ryder:1885:SPC

[Ryd85] John A. Ryder. On some of the protective contrivances developed by and in connection with the ova of various species of fishes. Transactions of the American Fisheries Society, 14(1):59–65, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ryder:1889:LLO

[Ryd89] John A. Ryder. The lateral line organs and the hyaline tissues of the head of the shad. Transactions of the American Fisheries Society, 18 (1):20–25, January 1889. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smiley:1882:CFG

[Smi82] Charles W. Smiley. Changes in the fisheries of the Great Lakes during the decade, 1870–1880. Transactions of the American Fisheries Society, 11(1):28–37, January 1882. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stearns:1885:GCP

[Ste85] Robt E. C. Stearns. The giant claims of Puget Sound. Transactions of the American Fisheries Society, 14(1):8–12, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stearns:1886:IUD

[Ste86] R. E. C. Stearns. Intentional and unintentional distribution of species. Transactions of the American Fisheries Society, 15(1):50–56, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stone:1884:APS

[Sto84] Livingston Stone. The artificial propagation of salmon in the Columbia River Basin. Transactions of the American Fisheries Society, 13(1):21– 31, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tomlin:1887:MLS

[Tom87] W. D. Tomlin. Migration of Lake Superior fish. Transactions of the American Fisheries Society, 16(1):60–64, January 1887. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

True:1885:PS

[Tru85a] F. W. True. Porpoise steak. Transactions of the American Fisheries Society, 14(1):38–40, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

True:1885:PFC

[Tru85b] Frederick W. True. The porpoise fishery of Cape Hatteras. Transactions of the American Fisheries Society, 14(1):32–38, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanCleef:1885:HRO

[Van85] J. S. Van Cleef. How to restore our trout streams. Transactions of the American Fisheries Society, 14(1):50–55, January 1885. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Washburn:1886:DSD

[Was86] F. L. Washburn. Deep sea dredging on the U.S. Steamer Albatross. Transactions of the American Fisheries Society, 15(1):17–21, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitaker:1886:MG

[Whi86] Herschel Whitaker. The Michigan grayling. Transactions of the American Fisheries Society, 15(1):59–67, January 1886. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Willcox:1884:FSF

[Wil84] Joseph Willcox. The Florida sponge fishery. Transactions of the American Fisheries Society, 13(1):67–109, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winslow:1884:PCF

[Win84] Francis Winslow. Present condition and future prospects of the oyster industry. Transactions of the American Fisheries Society, 13(1):148– 163, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Worth:1884:PSB

[Wor84] S. G. Worth. The propagation of the striped bass. Transactions of the American Fisheries Society, 13(1):209–212, January 1884. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Worth:1887:NCE

[Wor87] S. G. Worth. North Carolina encouragement to shellfish culture. Transactions of the American Fisheries Society, 16(1):53–54, January 1887. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).